# Building symmetries into data-driven manifold dynamics models for complex flows

### v1

@@ III Data-driven dimension reduction and dynamic modeling
In the following sections, we describe the steps involved in symmetry charting: 1) mapping data to the fundamental space, 2) finding a minimal-dimensional coordinate representation, and 3) evolving the minimal-dimensional state and symmetry indicators forward through time. The parts of this procedure separating it from other data-driven reduced order models are how we map to the fundamental and how we evolve the symmetry indicators through time. Our approach guarantees equivariance under symmetry operations.
@@ III.1 Map to fundamental domain
First we discuss a method to map trajectories of Kolmogorov flow to the fundamental domain where symmetries are factored out [30]. The first step is to identify a set of indicators that are related to the group of symmetries. In order to find these indicators we consider the effect of these symmetry operations in Fourier space. After Fourier transforming the equations and simplifying these are the actions of the symmetry operations on the Fourier coefficients:
@@
$\mathscr{T}_{l}:\hat{\omega}(k_{x},k_{y})\rightarrow\hat{\omega}(k_{x},k_{y})e^{-ik_{x}l},$ (9)
where $\bar{\cdot}$ denotes the complex conjugate.
Now, we will show how these symmetry operations act on the Fourier coefficients in specific ways that we can exploit to classify any given state as lying within a specific region or domain of state space. We refer to this classification of the continuous symmetry as the “phase” and the discrete symmetry as the “indicator”. To compute the phase and remove the continuous symmetry $\mathscr{T}_{l}$, we use the First Fourier mode method-of-slices [31, 32]. This method involves computing the spatial phase $\phi_{x}$ of the $k_{x}=1$ and $k_{y}=0$ Fourier mode: $\phi_{x}(t)=\operatorname{atan}2\left\{\hat{\omega}_{I}(1,0),\hat{\omega}_{R}(1,0)\right\}$. Then, to phase-align the data, we shift the vorticity snapshots such that this mode is a pure cosine: $\omega_{l}(x,y,t)=\mathcal{F}^{-1}\left\{\mathcal{F}\{\omega(x,y,t)\}e^{-ik\phi_{x}(t)}\right\}$. Doing this ensures that the snapshots lie in a reference frame where no translation happens in the $x$ direction. From Eq. 9 we see that this operation only modifies Fourier coefficients with $k_{x}\neq 0$. Thus, if we use $k_{x}=0$ coefficients to determine the discrete indicator it will be independent of the phase.
For the shift-reflect operation, we consider the $(k_{x}=0,k_{y}=1)$ mode. In Eq. 7 the shift-reflect operation modifies this Fourier coefficient in this manner: $\hat{\omega}_{R}(0,1)+i\hat{\omega}_{I}(0,1)\rightarrow\hat{\omega}_{I}(0,1)-i\hat{\omega}_{R}(0,1)$. This shows that the shift-reflect operation maps between different quadrants of the $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ plane. Next, we consider the rotation operation on the $(k_{x}=0,k_{y}=2)$ mode. In Eq. 8 we see this operation simply flips the sign of the imaginary component $\hat{\omega}_{I}(0,2)\rightarrow-\hat{\omega}_{I}(0,2)$. Thus, applying the eight possible discrete symmetry operations (including the identity operation) to a vorticity field will map us to eight unique points in the $\hat{\omega}_{R}(0,1)$, $\hat{\omega}_{I}(0,1)$, $\hat{\omega}_{I}(0,2)$ space. As such, we may classify each octant with an $\mathcal{I}$ between $0-7$. This means that any snapshot can be mapped to a fundamental domain and be uniquely identified by $\mathcal{I}$. In Fig. 4 we show the indicators for $\text{Re}=13.5$. The colored curves correspond to the sections of the different RPOs that are related by discrete symmetries. Notice that each section of the different RPOs can be uniquely identified by $\mathcal{I}$. Table 1 shows how any snapshot in any octant can be mapped to the $\mathcal{I}=0$ octant by applying the corresponding discrete symmetry operations. These same indicators extend to the chaotic case $\text{Re}=14.4$ and in the rest of this work we refer to the fundamental domain as the space corresponding to $\mathcal{I}=0$. Fig. 5 shows the state-space projection of a chaotic trajectory at $\text{Re}=14.4$ into the region $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)-\hat{\omega}_{I}(0,2)$, where the different colors give the indicator for the chart in which that data point lies. More generally, we note that any point in the state space of the system can represented by a point in the fundamental domain along with an indicator $\mathcal{I}$ and phase $\phi_{x}$.
This representation is satisfactory for static data, but not for modeling trajectories represented in the fundamental domain, because when a trajectory exits one region, its corresponding trajectory in the fundamental domain leaves there and instantaneously enters at a different point: i.e. the time evolution is discontinuous in this representation.
To address this issue we use the concept of charts and atlases that is fundamental to the study of manifolds [14], in a way that is closely related to that of Floryan & Graham [15]. An atlas for a manifold is a collection of patches, each of which must overlap with its neighbors, that are called charts. In each chart, a local coordinate representation can be found, and for each pair of overlapping charts, a transition map takes coordinates in one chart to those in the neighboring one – with this coordinate tranformation there are no discontinuities. To use this formalism in the present case requires us to expand the fundamental domain so that it overlaps with its neighbors, so the fundamental chart is simply the fundamental domain, comprised of “interior points” plus the overlap region (“exterior points”), and there are seven other charts, each generated by the symmetry operation acting on the fundamental chart including the overlap regions. We choose a coordinate representation for all charts that corresponds to that for the fundamental chart, and for a trajectory that leaves the fundamental domain with e.g. $\mathcal{I}=4$ and reenters with $\mathcal{I}=6$, the transition map is the function that moves the point in the fundamental chart from the “exit” of the $\mathcal{I}=4$ chart to the “entrance” of the $\mathcal{I}=6$ chart.
In Fig. 4(a) we show the state space projection of a trajectory, identifying interior points with solid markers and exterior points with open markers. We isolate the region close to an RPO in Fig. 4(b) to make this more clear. Solid markers correspond to data points that lie in the interior of a chart. Open markers correpond to the data points in the connecting chart that will be included together with the interior data. Many different approaches can be taken to identify points to include in the exterior. Here we simply identified all points in a small volume around each octant as exterior points. We did this by finding the dimensions of a box surrounding the octant (by calculating the maximum absolute value in the $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)-\hat{\omega}_{I}(0,2)$ directions) and then increasing the size of this box by expanding each direction $20\%$. Any point that lies in the volume between the octant and this expanded domain is an exterior point. We varied the percent increase to $5\%$, $10\%$, and $30\%$ and observed no change in the results.
Now that we have identified these exterior points we can continue with the procedure of mapping the data to the fundamental chart. To do this, we apply the appropriate discrete operations to the interior and exterior points of each chart to map the state such that the interior points fall in $\mathcal{I}=0$ ($\hat{\omega}_{R}(0,1)>0$, $\hat{\omega}_{I}(0,1)>0$, $\hat{\omega}_{I}(0,2)>0$). Fig. 5(a) shows the colored clusters mapped onto the positive octant, dark blue cluster. Figures 5(b) and 5(c) show projections in the planes $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ and $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,2)$ respectively of data mapped to the positive octant. The dashed lines correspond to the boundaries between the different octants. We see clearly that mapping the data to the fundamental chart leads to a much higher density of data points and thus to better sampling than if we considered the whole state space without accounting for the symmetries. Following the mapping to the fundamental chart we phase-align the snapshots and refer to this data as $\tilde{\omega}$ from this point on.
Figure 4: State-space projection of the different RPOs into the plane $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)-\hat{\omega}_{I}(0,2)$ for $\mathrm{Re}=13.5$. The different colors are for different discrete symmetry indicators which are shown in the legend.
@@
Table 1: Indicators $\mathcal{I}$ and the corresponding operations required to map to $\mathcal{I}=0$
@@
 $\begin{array}[]{cc}\hline\cr\hline\cr\mathcal{I}&\text{\; \; \; \; Discrete operations }\\ \hline\cr 0&\text{None}\\ 1&\mathscr{R}\omega\\ 2&\mathscr{S}^{3}\omega\\ 3&\mathscr{S}^{3}\mathscr{R}\omega\\ 4&\mathscr{S}^{2}\mathscr{R}\omega\\ 5&\mathscr{S}\mathscr{R}\omega\\ 6&\mathscr{S}^{2}\omega\\ 7&\mathscr{S}\omega\\ \hline\cr\hline\cr\end{array}$
@@
Figure 5: (a) State-space projection of a trajectory into the subspace $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)-\hat{\omega}_{I}(0,2)$ for $\mathrm{Re}=14.4$. Colors correspond to the discrete symmetry indicator. (b) Zoom into the region near the RPO corresponding to $\mathcal{I}=1,7$. Solid markers correspond to interior points and open markers to exterior points – e.g.  the blue solid markers within an open orange marker are in the interior region of chart 7 and exterior region of chart 1.
@@
Figure 6: (a) Snapshots mapped to the fundamental chart ($\hat{\omega}_{R}(0,1)>0$, $\hat{\omega}_{I}(0,1)>0$, $\hat{\omega}_{I}(0,2)>0$). (b) and (c) are the state-space projection of the trajectory into the planes $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ and $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,2)$, respectively. The dotted lines give the boundary between interior and exterior points.
@@ III.2 Finding a manifold coordinate representation with IRMAE-WD
In the previous section we discussed how to obtain the fundamental chart. We have defined a natural subdivision of the state space, where the invariant manifold of the dynamics is represented by an atlas of charts that are related by symmetry operations. This means that we do not need to learn the full invariant manifold, only the piece in the fundamental chart. To find a low-dimensional representation of the fundamental chart we use a variant of a common machine-learning architecture known as an undercomplete autoencoder (AE), whose purpose is to learn a reduced representation of the state such that the reconstruction error with respect to the true data is minimized. We flatten the vorticity field $\tilde{\omega}$ such that $N=32\times 32=1024$. The AE consists of an encoder and a decoder. The encoder maps from the full space $\tilde{\omega}\in\mathbb{R}^{N}$ to the lower-dimensional latent space $h(t)\in\mathbb{R}^{d_{h}}$, where ideally $d_{h}=d_{\mathcal{M}}\ll N$, and the decoder maps back to the full space $\tilde{\omega}_{r}$.
@@
Fig. 7a shows the IRMAE-WD framework. The encoder, denoted by $E\left(\tilde{\omega};\theta_{E}\right)$ reduces the dimension from $N$ to $d_{z}$. We then include a linear network $\mathcal{W}\left(\cdot;\theta_{W}\right)$ between the encoder and the decoder which consists of several linear layers (matrix multiplications). Finally, the decoder $\tilde{\omega}_{r}=D\left(z;\theta_{D}\right)$ maps back to the full space. An $L_{2}$ (“weight decay”) regularization to the weights is also added, with prefactor $\lambda$. The loss function for this architecture is
@@
After training, we can perform SVD on the covariance matrix of the encoded data matrix $Z$ to obtain the singular vectors $U$ and singular values $S$ as shown in Fig. 7b. Projecting $z$ onto $\hat{U}^{T}$ gives an orthogonal representation $h^{+}=U^{T}z$ as illustrated in Fig. 7c. Then, by choosing only the singular values above some very small threshold (typically $\gtrsim 6$ orders of magnitude smaller than the leading singular values), we may project down to fewer dimensions by projecting onto the corresponding singular vectors $U$, denoted $\hat{U}$ to yield the low-rank manifold representation $h=\hat{U}z$ (Fig. 7d). We refer to Table 2 for details on the architecture.

### v2

@@ III Data-driven dimension reduction and dynamic modeling
In the following subsections, we describe the steps involved in symmetry charting: 1) developing criteria to define a fundamental chart in the the original (high-dimensional) state space and what symmetry operations are need to map any point in state space to that chart, 2) finding a minimal-dimensional coordinate representation, i.e. the fundamental chart for the invariant manifold, and 3) evolving both the minimal-dimensional state and the symmetry indicators forward through time. The parts of this procedure separating it from other data-driven reduced-order models are how we map to the fundamental domain and how we evolve the symmetry indicators through time. Our approach guarantees that the reduced-dimensional model maintains the equivariance properties of the original dynamical system.
@@ III.1 Map to fundamental domain
In the example shown in Fig. 1, only a simple, discrete symmetry was present, and it was straightforward to identify an indicator value $\mathcal{I}$ corresponding with each symmetry operation that would map a general state point $\boldsymbol{y}$ to the fundamental domain. This subsection illustrates how to do so for the symmetries of the Kolmogorov flow. To illustrate the generality of the approach, in Appendix B, we outline the same process for plane channel flow. As noted above, for the Kolmogorov flow we consider, there are eight distinct discrete symmetry operations, so we define an indicator $\mathcal{I}\in\left\{0,1,2,\ldots,7\right\}$ as well as the continuous translation in $x$, which we treat as a spatial phase shift $\phi_{x}$. We describe here how to take an arbitrary vorticity field $\omega(x,y)$ and rewrite it uniquely as the set $\left\{\omega_{F}(x,y),\phi_{x},\mathcal{I}\right\}$, where $\omega_{F}(x,y)$ is a state point in a chosen fundamental domain, i.e. with the discrete and continuous symmetries factored out. Now a trajectory $\omega(x,y,t)$ can be represented as $\left\{\omega_{F}(x,y,t),\phi_{x}(t),\mathcal{I}(t)\right\}$.
The criteria that determine $\phi_{x}$ and the various indicator indices $\mathcal{I}$ are most conveniently represented in Fourier space. With $x$- and $y$-wavenumbers $k_{x}$ and $k_{y}$, respectively, operations (3)-(5) applied to the vorticity field at a particular time instant become
@@
$\mathscr{T}_{l}:\hat{\omega}(k_{x},k_{y})\rightarrow\hat{\omega}(k_{x},k_{y})e^{ik_{x}l},$ (9)
where $\hat{\cdot}$ denotes Fourier transform $\mathcal{F}$ and $\bar{\cdot}$ denotes complex conjugate. Moreover, since the vorticity field is real, $\hat{\omega}(k_{x},-k_{y})=\bar{\hat{\omega}}(-k_{x},k_{y})$ and $\hat{\omega}(-k_{x},-k_{y})=\bar{\hat{\omega}}(k_{x},k_{y})$.
To compute the phase $\phi_{x}$ for a vorticity field, we use the “First Fourier mode method-of-slices” [35, 36]. This method involves computing the spatial phase $\phi_{x}$ of the Fourier mode $k_{x}=1,k_{y}=0$: using standard trigonometric identities, this mode can be written as $2||\hat{\omega}(1,0)||\cos(x+\phi_{x})$, with $\phi_{x}(t)=\operatorname{atan}2\left\{\hat{\omega}_{I}(1,0),\hat{\omega}_{R}(1,0)\right\}$, where subscripts $R$ and $I$ indicate real and imaginary part. Then, to factor out the translation, we shift the vorticity snapshots such that this mode is a pure cosine: $\omega_{\phi_{x}}(x,y,t)=\mathcal{F}^{-1}\left\{\mathcal{F}\{\omega(x,y,t)\}e^{-ik_{x}\phi_{x}(t)}\right\}$ The translation of the solution in the $x$-direction is captured by the time-dependence of $\phi_{x}$.
Now we turn to the discrete symmetries. As noted above, for a given state $\omega(x,y)$ there are seven other states related by the symmetry operations. Our aim is thus to identify eight regions of state space that are related to one another by these operations. We will define these regions using simple sign criteria based on how the dominant (low wavenumber) Fourier modes transform under $\mathscr{S}$ and $\mathscr{R}$. We will need three such criteria to identify the eight regions in state space. To simplify the analysis, we use coefficients with $k_{x}=0$ because these values are invariant to translations in $x$ and thus unchanged when the translation symmetry is factored out.
The actions of $\mathscr{S}$ and $\mathscr{R}$ on $\hat{\omega}(0,1)$ and $\hat{\omega}(0,2)$ are as follows:
$\mathscr{S}:\hat{\omega}_{R}(0,1)+i\hat{\omega}_{I}(0,1)\rightarrow\hat{\omega}_{I}(0,1)-i\hat{\omega}_{R}(0,1),$ (10)
$\mathscr{S}:\hat{\omega}_{R}(0,2)+i\hat{\omega}_{I}(0,2)\rightarrow\hat{\omega}_{R}(0,2)+i\hat{\omega}_{I}(0,2),$ (11)
$\mathscr{R}:\hat{\omega}_{R}(0,1)+i\hat{\omega}_{I}(0,1)\rightarrow\hat{\omega}_{R}(0,1)-i\hat{\omega}_{I}(0,1),$ (12)
$\mathscr{R}:\hat{\omega}_{R}(0,2)+i\hat{\omega}_{I}(0,2)\rightarrow\hat{\omega}_{R}(0,2)-i\hat{\omega}_{I}(0,2).$ (13)
Observe that the shift-reflect operator $\mathscr{S}$ modifies the $(k_{x}=0,k_{y}=1)$, but not the $(k_{x}=0,k_{y}=2)$ coefficients. I.e. $\mathscr{S}$ maps between different quadrants of the $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ plane: $\mathscr{S}^{3},\mathscr{S}^{2}$, and $\mathscr{S}$ map from quadrants 2, 3, and 4, of this plane to the first quadrant, respectively. The rotation operator $\mathscr{R}$ modifies both the $(k_{x}=0,k_{y}=1)$ and $(k_{x}=0,k_{y}=2)$ coefficients by flipping the sign of the imaginary component. We now can take the fundamental domain to be the set of points whose projection into the $(\hat{\omega}_{R}(0,1)$, $\hat{\omega}_{I}(0,1)$, $\hat{\omega}_{I}(0,2))$ lies in the (+,+,+)(+,+,+)( + , + , + ) octant, and use the above observations to determine what operation maps a point whose projection lies in any other octant into the first. Table 1 shows these discrete operations, identifying each octant by the signs of the Fourier coefficients. For example, for a data point in the (,,)(-,-,-)( - , - , - ) octant, the operation $\mathscr{R}\mathscr{S}^{3}$ maps it into the fundamental domain. The entries for $\mathscr{R},\mathscr{S},\mathscr{S}^{2}$ and $\mathscr{S}^{3}$ follow directly from the discussion above, with the remaining entries obtained by inspection.
In summary, for any vorticity field, the signs of $\hat{\omega}_{R}(0,1)$, $\hat{\omega}_{I}(0,1)$, $\hat{\omega}_{I}(0,2)$ specify an octant, and one of the eight symmetry operations will bring the field into the first (+,+,+)(+,+,+)( + , + , + ) octant. Now we can fully characterize the fundamental domain as the region in state space where $\phi_{x}=0$ and $\mathcal{I}=0$. For any vorticity field $\omega(x,y,t)$ we can determine and factor out $\phi_{x}(t)$ and $\mathcal{I}(t)$ using the above criteria to give a point $\omega_{F}(x,y,t)$.
@@
Table 1: Indicators $\mathcal{I}$, corresponding operations required to map to $\mathcal{I}=0$, the fundamental domain, and signs of the relevant Fourier amplitudes.
@@
 $\begin{array}[]{ccccc}\hline\cr\hline\cr\mathcal{I}&\text{\; \; \; \; Discreteoperation }&\textrm{sgn}(\hat{\omega}_{R}(0,1))&\textrm{sgn}(\hat{\omega}_{I}(0,1))&\textrm{sgn}(\hat{\omega}_{I}(0,2))\\ \hline\cr 0&\text{Identity}&+&+&+\\ 1&\mathscr{R}&+&{\color[rgb]{0,0,0}-}&{\color[rgb]{0,0,0}-}\\ 2&\mathscr{S}&-&+&+\\ 3&\mathscr{R}\mathscr{S}^{3}&-&{\color[rgb]{0,0,0}-}&-\\ 4&\mathscr{R}\mathscr{S}^{2}&-&{\color[rgb]{0,0,0}+}&-\\ 5&\mathscr{R}\mathscr{S}&+&{\color[rgb]{0,0,0}+}&-\\ 6&\mathscr{S}^{2}&-&-&+\\ 7&\mathscr{S}^{3}&+&-&+\\ \hline\cr\hline\cr\end{array}$
Fig. 4 shows the state-space projection of four symmetry-related relative periodic orbits (RPOs) at $\mathrm{Re}=13.5$ into the subspace $(\hat{\omega}_{R}(0,1),\hat{\omega}_{I}(0,1),\hat{\omega}_{I}(0,2))$, where the different colors give the indicator for the octant in which that data point lies. Notice that each section of the different RPOs can be uniquely identified by $\mathcal{I}$. Fig. 5(a) shows the same, for a chaotic trajectory at $\text{Re}=14.4$.
As we see in Figures 4 and 5(a), a trajectory of the dynamics may pass through several or even all of these octants. To handle the transitions between these regions, we use the concept of charts and atlases that is fundamental to the study of manifolds [14], in a way that is closely related to that of Floryan & Graham [15]. An atlas for a manifold is a collection of regions of the manifold, each of which must overlap with its neighbors, that are called charts. In each chart, a local coordinate representation can be found, and for each pair of overlapping charts, a transition map takes coordinates in one chart to those in the neighboring one – with this coordinate transformation there are no discontinuities. Note that one can define an atlas of charts on both the ambient state space $\mathbb{R}^{N}$ and on the invariant manifold $\mathcal{M}$ of the dynamics. (The phase variable can be taken to be an additional coordinate for either the ambient state space or the invariant manifold.) The fundamental chart for $\mathcal{M}$ is just the portion of $\mathcal{M}$ in the fundamental chart of the ambient space, and it is related to the other charts via the indicators $\mathcal{I}$. In this section we focus on the ambient space.
To construct the fundamental chart, we expand the fundamental domain so that it overlaps with neighboring regions of state space. That is, the fundamental chart is simply the fundamental domain, comprised of “interior points”, whose projection into the $(\hat{\omega}_{R}(0,1),\hat{\omega}_{I}(0,1),\hat{\omega}_{I}(0,2))$ subspace lie in the first octant, plus an overlap region containing “exterior points” whose projections lie just outside the first octant. Specifically, for a given data set, we find the point whose projection into the $(\hat{\omega}_{R}(0,1),\hat{\omega}_{I}(0,1),\hat{\omega}_{I}(0,2))$ subspace has largest magnitude, i.e. that maximizes $\hat{\omega}_{R}(0,1)^{2}+\hat{\omega}_{I}(0,1)^{2}+\hat{\omega}_{I}(0,2)^{2}$. Denoting the coordinates of this projected point as $a,b,c$, we then define a box in the first octant with one corner at the origin and the far corner at $(\left|a\right|,\left|b\right|,\left|c\right|)$. That will be the fundamental domain. To make the fundamental chart, we simply expand that box in all directions by $20\%$ so it overlaps into all seven other octants. Any point whose projection lies in the volume between the original (first octant) box and this expanded box is an exterior point. We varied the percent increase to $5\%$, $10\%$, and $30\%$ and observed no change in the results. Expanding the other seven octants in the same manner generates eight overlapping charts. We choose a coordinate representation for all charts that corresponds to that for the fundamental chart, and for a trajectory that leaves the fundamental domain with e.g. $\mathcal{I}=4$ and reenters with $\mathcal{I}=6$, the transition map is the function that moves the point in the fundamental chart from the “exit” of the $\mathcal{I}=4$ chart to the “entrance” of the $\mathcal{I}=6$ chart.
@@
In Fig. 5 we show the state space projection of a chaotic trajectory at $\mathrm{Re}=14.4.$ Points surrounded by open circles are interior points for one chart and exterior points to a neighboring one. E.g. the blue solid markers within an open orange marker are in the interior region of chart 7 and exterior region of chart 1 and vice versa.
Now that we have identified these exterior points we can map the entire data set to the fundamental chart. To do this, we apply the appropriate discrete operations to the interior and exterior points of each chart to map the state such that the interior points fall in the first octant $\mathcal{I}=0$ ($\hat{\omega}_{R}(0,1)>0$, $\hat{\omega}_{I}(0,1)>0$, $\hat{\omega}_{I}(0,2)>0$). For the chaotic flow at $\mathrm{Re}=14.4$, Fig. 6(a) shows an entire trajectory mapped into the fundamental chart, with the colors indicating the indicator values for each point. Figures 6(b) and 6(c) show projections in the planes $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ and $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,2)$, respectively, of the mapped data. The dashed lines correspond to the boundaries between the different octants. Observe that mapping the data to the fundamental chart leads to a much higher density of data points and thus to better sampling than if we considered the whole state space without accounting for the symmetries. From here on, we refer to data has been phase-aligned and mapped into the fundamental chart as $\tilde{\omega}$.
Figure 4: State-space projection of the different RPOs at $\mathrm{Re}=13.5$ into the subspace $(\hat{\omega}_{R}(0,1),\hat{\omega}_{I}(0,1),\hat{\omega}_{I}(0,2))$. The colors indicate the discrete symmetry operations (Table 1) required to map the data to the fundamental domain.
Figure 5: (a) Projection of a chaotic trajectory at $\mathrm{Re}=14.4$ into the subspace $(\hat{\omega}_{R}(0,1),\hat{\omega}_{I}(0,1),\hat{\omega}_{I}(0,2))$. The colors indicate the discrete symmetry operations (Table 1) required to map the data to the fundamental domain. (b) Zoom into the region near the RPO corresponding to $\mathcal{I}=1,7$. Colors of solid markers correspond to which octant they occupy. Points surrounded by open circles are interior points for one chart and exterior points to a neighboring one. E.g.  the blue solid markers within an open orange marker are in the interior region of chart 7 and exterior region of chart 1 and vice versa.
@@
Figure 6: (a) Snapshots mapped to the fundamental chart ($\hat{\omega}_{R}(0,1)>0$, $\hat{\omega}_{I}(0,1)>0$, $\hat{\omega}_{I}(0,2)>0$) in the ambient space. (b) and (c) are the state-space projection of the trajectory into the planes $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,1)$ and $\hat{\omega}_{R}(0,1)-\hat{\omega}_{I}(0,2)$, respectively. The dotted lines give the boundary between interior and exterior points.
@@ III.2 Finding a manifold coordinate representation with IRMAE-WD
In the previous section we demonstrated how to obtain the fundamental chart in the ambient space. Now we turn to representing the lower-dimensional invariant manifold $\mathcal{M}$, noting that we do not need to represent the entire manifold but only the portion that is in the fundamental chart of the ambient space, i.e. the region illustrated in Fig. 6 . This will be the fundamental chart for $\mathcal{M}$, with the other charts related to it via the symmetry operations. Mapping all the data in the ambient space into the fundamental chart thus represents an eightfold augmentation of the data set that is available to determine the coordinate representation for the fundamental chart on $\mathcal{M}$ (i.e. in the autoencoder latent space).
To find a low-dimensional representation of the fundamental chart on $\mathcal{M}$ we use a variant of a common machine-learning architecture known as an undercomplete autoencoder (AE), whose purpose is to learn a reduced representation of the state such that the reconstruction error with respect to the true data is minimized. We flatten the vorticity field $\tilde{\omega}$ such that $N=32\times 32=1024$. The AE consists of an encoder and a decoder. The encoder maps from the full space $\tilde{\omega}\in\mathbb{R}^{N}$ to the lower-dimensional latent space $h(t)\in\mathbb{R}^{d_{h}}$, where ideally $d_{h}=d_{\mathcal{M}}\ll N$, and the decoder maps back to the full space $\tilde{\omega}_{r}$. The vector $h(t)$ is a set of coordinates for representing the fundamental chart on $\mathcal{M}$.
@@
Fig. 7a shows the IRMAE-WD framework. The encoder, denoted by $z=E\left(\tilde{\omega};\theta_{E}\right)$ reduces the dimension from $N$ to the bottlneck dimension $d_{z}$. We then include a linear network $\mathcal{W}\left(\cdot;\theta_{W}\right)$ between the encoder and the decoder that consists of several linear layers (matrix multiplications). Finally, the decoder $\tilde{\omega}_{r}=D\left(z;\theta_{D}\right)$ maps back to the full space. An $L_{2}$ (“weight decay”) regularization to the weights is also added, with prefactor $\lambda$. The loss function for this architecture is
@@
After training, we perform SVD on the covariance matrix of the encoded data matrix $Z$ to obtain the singular vectors (columns of $U$) and singular values (diagonal elements $\sigma_{i}$) of $S$ as shown in Fig. 7b. Projecting $z$ onto $\hat{U}^{T}$ gives an orthogonal representation $h^{+}=U^{T}z$ as illustrated in Fig. 7c. Then, by choosing only the singular values $\sigma_{i}$ above some very small threshold (typically $\gtrsim 6$ orders of magnitude smaller than the leading singular values), we may project down to fewer dimensions by projecting onto the corresponding singular vectors, contained in the matrix $\hat{U}$, to yield the low-rank manifold representation $h=\hat{U}z$ (Fig. 7d). We refer to Table 2 for details on the architecture.